OpenStudy (anonymous):

Joe wants to buy a new tennis racquet that costs $200 from the factory the sports shop marks up the price by 60% Joe bought the racquet when the shop had 25% off sale how much profit did the shop make

OpenStudy (anonymous):

racket cost \[200\times 1.6\times .75 = 200\times 1.2 =\$ 240\]

OpenStudy (anonymous):

((200*1.6) 0.75) - 200 = 40 Joe had to pay 75% of the retail price in cash, $240. The store paid $200 and the difference is $40.